# Electron identification and hadron discrimination using Cherenkov radiation in air and SiPMs

### v1

@@ 1 Introduction
In [SiPM1, SiPM2, SiPM3, SiPM4] it was quantitatively shown that the Cherenkov radiation produced at the passage of a charged particle through the standard, factory-supplied protection layer of the SiPM itself can produce very high number of firing cells (SPADs with area tens of microns square). Considering a refractive index of 1.56 and a thickness of few hundreds micron for the Silicone material of the protection layer, the Cherenkov angle reaches $\approx$50 degrees, covering a quite large area of few mm2. In these conditions it is possible to directly detect charged particles and reach very good efficiency and time resolution around 20 ps [SiPM3, SiPM4]. These results opened the possibility to use just a single SiPM as an optimal detector for Time Of Flight (TOF) measurements.
The results reported in the above papers also indicate a very reduced number of firing SPADs when no protection layer is present. For such sensors, events with more than one SPAD are mainly due to cross-talk processes [SiPM3, SiPM4].
@@
To test the above possibility, in this paper data collected at T10 test beam at CERN with momenta from 1.5 to 10 GeV/$c$ are used. The apparatus is the same described in [SiPM3]. At lower momenta the fraction of electrons/positrons in the beam increases thus allowing to compare the SiPM (without protection) signal with the one observed for pions and protons. In Section 2 the experimental set-up and the analysis methods are described. The results are reported in Section 3 and will be compared with a simple simulation in Section 4. In Section 5 a possible optimization of the parameters involved in such study (length of air traversed, different traversed gas, different SiPM technology) is discussed.
@@ 2.2 Beam test setup
The SiPMs were tested at the CERN PS T10 particle beam in the momentum range 1.5-10 GeV/$c$. The beam composition, evaluated via simulation and data, is reported in [vandijk]: while at high momenta the beam is mainly made of protons and $\pi^{+}$, at lower energies the fraction of positrons increases and becomes dominant. In the following, we will neglect the charge, which is uninfluential for our purposes, and use generically the terms pions and electrons.
@@ 2.3 Analysis method
The analyzed data were collected with a beam momentum of 1.5 GeV/$c$ (a second sample at 10 GeV/c was also collected to check the simulation). The reason for this choice is twofold: first, it enhances the electron component in the beam, which is composed of approximately 60% electrons, 28% pions, and 12% protons [vandijk]. Second, it enables the separation of electrons and pions from protons by exploiting the time of flight between the signals from the two LGAD detectors placed at the ends of the telescope, as visible in Fig. 1 left (more details in [SiPM3]). In the figure a q-Gaussian fit [q-gauss] was applied on both distributions.
Figure 1: Left panel: time of flight of the beam particles computed as the time difference between the two LGADs in the telescope at 1.5 GeV/$c$. The fits are performed with two q-Gaussian, the left one relative to the proton content. Right panel: LGAD time sum versus time difference distribution.
At this beam momentum, electrons are the only particles that produce Cherenkov radiation in the air gap between sensors in the telescope. Electrons and pions cannot be distinguished using the LGADs’ time information; however, their charge spectra can still be determined. The proton charge spectrum is used as a proxy for pions normalized to the expected pion fraction [vandijk], since neither pions nor protons are able to generate Cherenkov radiation in air.
The protons sample has been selected with rectangular cuts on the LGAD time sum versus time difference distribution (Fig. 1 right) as 0.2 $t_{LGAD1}-t_{LGAD2}<$ 0.4 ns and -1.2 $t_{LGAD1}+t_{LGAD2}< -$0.85 ns while for electrons and pions the cuts are 0 $t_{LGAD1}-t_{LGAD2}<$ 0.16 ns and -1.2 $t_{LGAD1}+t_{LGAD2}< -$0.7 ns. With these cuts the fraction of protons remaining in the electron/pion sample is below 1%, while the electron/pion contamination in the proton sample is $\approx$ 4-5%.
@@ 3 Results
The measured amplitudes for the proton sample is reported in Fig. 2 top-left: as expected, it is peaked at one SPAD ($\sim$31 mV), with the entries at 0 mV (0 SPAD) compatible with the value of the fill factor of the sensor and with peaks at 2 and 3 SPADs due to crosstalk. The amplitude distribution for pions is expected to be identical to that of protons. The electron distribution is then obtained by taking the amplitude distribution of events identified as electron/pion candidates and subtracting the proton distribution, normalized to the expected pion fraction of 32%[vandijk]. The pion/electron spectrum is reported in Fig. 2 top-right with the estimated fraction of pions and electrons. The electron spectrum, after the pion subtraction, is instead reported in Fig. 2 bottom.
@@ 4 Comparison with simulation
The results reported in the previous section rely mainly on the purity of the selected proton sample and on an accurate knowledge of the actual composition of the particle beam. To estimate the impact of these experimental uncertainties the results were cross-checked against the predictions of a toy Monte Carlo (MC) simulation to assess their robustness. The simulation works as follows: two hits are randomly generated on the two LGADs and the track is then propagated to the SiPM. The number of Cherenkov photons created in the air volume upstream the sensor has been calculated based on the Frank-Tamm formula [FrankTamm]. The detection efficiency for the charged particle was assumed equal to the fill factor according to the manufacturer’s specifications, as well as the crosstalk and the PDE as a function of the photon wavelength and SiPM overvoltage. The amplitude distribution is obtained by multiplying the number of fired cells by the average amplitude of 1 SPAD signal assumed as 31 mV according to the acquired data. A gaussian spread of 3.3 mV was applied driven by data observation. Given the binary nature of the SPAD response, multiple photons hitting the same pixel are counted as one. A still-not-fully-satisfactory modeling of the noise is applied to the simulated data based on the measured RMS of the baseline which, however, does not appear to have a significant effect on data reproducibility.
@@
A qualitative comparison between the measured spectra and MC predictions is shown in Fig. 3. In the left pad, the measured amplitude spectrum for protons with a momentum of 1.5 GeV/$c$ is compared with simulations. Here and in the following, the two histograms under comparison are normalized to the same number of entries. The measured spectrum shows a small peak at 3 SPADs which is not predicted by the toy MC. This discrepancy ($\sim 3\%$ of events with more than 2 SPADs) mainly arise from a small electron contamination in the proton sample selected by LGAD timing information, $\sim 2-3\%$; the remaining part may originate from events in which crosstalk affects more than one pixel, which is not included in the MC, or from particle near the border of two SPADs. In the same figure the right panel shows a comparison between measured amplitude spectra and simulations at a beam momentum of 10 GeV/$c$. At this momentum, the beam consists mainly of pions (30%) and protons (70%)[vandijk], with pions producing Cherenkov radiation in air, while protons remain below threshold. In this case, the agreement between the two distributions is qualitatively good and the problems noticed in the previous plot are less evident due to the higher number of SPADs firing. Finally, Fig. 4 compares the measured and simulated amplitude spectra for electrons at 1.5 GeV/$c$ : here too, the comparison shows good agreement.
@@ 5 A simple optimization of the key parameters
Before using the toy MC to obtain an approximate estimate of the potential performance of such a system, a number of parameters were optimized. It was observed that the SiPM coverage has a significant impact, with larger covered areas able to collect more photons produced by the incident charged particle. No significant changes in performance were observed reducing the cell size, within the range tested. A comparison was therefore made between SiPMs with active areas of 3.20 $\times$ 3.12 mm2, as the one used in the test beam data, and a hypothetical 6 $\times$ 6 mm2, both featuring square SPADs with a pitch of 40 µm. To determine the optimal thickness of the air layer in front of the sensor, the simulation considers particles impinging the SiPM in the center. The over-voltage was set to 6 V.
@@
The optimization was carried out using the plot shown in Fig. 5, where the electron efficiency was calculated considering particles with 1.5 GeV/$c$ of momentum and by selecting only events in which three or more SPADs were fired. As can be seen, while below 4 cm the curves relative to the two different size SiPM overlap, at higher distances the larger-area SiPM exhibits a significantly higher efficiency and this is due to the larger surface area, which allows more Cherenkov photons to be collected. The figure also indicates that for the 6 $\times$ 6 mm2 sensor beyond 15 cm of air no appreciable variations in the number of collected photons are observed due to the limited size of the simulated sensor. In the following simulations, the response of the larger SiPM will be considered using a distance of 15 cm of traversed air. Notice that the goal of this paper is to investigate the method to identify particles via the SiPM response and there is no attempt to optimize the detector configuration as it should be done for specific applications.
Since electron identification is performed by considering signals with an amplitude higher than the cross-talk value, it is clear that this value represents a fundamental parameter. Recently, the FBK NUV-HD SiPM technology has been upgraded with metal-filled deep trench isolation (NUV-HD-MT [NUV-HD-MT]) between the SiPM cells, which enables an almost complete suppression of internal optical cross-talk. In the simulation, therefore, in order to select the optimal configuration for the proposed detector, the cross-talk values measured by FBK for this new technology were used. The PDE also impacts performance, as it parameterizes the probability of detecting the produced Cherenkov photons. Assuming the validity of the simulation also at overvoltages higher than 2 V, a value of 6 V was chosen to profit of the higher PDE. The cross-talk and PDE values corresponding to OV = 6 V were taken from [NUV-HD-MT].
@@
Figure 5: Electron detection efficiency versus the distance in air traversed by the particle. Electrons of 1.5 GeV/$c$ and two different SiPMs surfaces are considered. The Toy- MonteCarlo parameters used are: NUV-HD-MT, OV = 6 V, Thr = 3 SPAD.
@@
Starting from the excellent timing capability of SiPM with a protection layer reported in [SiPM3], the approach discussed in the present paper suggests that configurations with two consecutive SiPMs, one without and one with protection layer, with an air gap in front of the first one, could be used to simultaneously support electron identification and TOF measurements, potentially covering PID and particle rejection in different momentum ranges.
@@ 6 Conclusions
Together with the already published results on timing performances of SiPMs with a protection layer [SiPM3, SiPM4], the present results underline that the Cherenkov light produced in the simplest radiators in front of the SiPM, i.e. the factory-supplied silicone protective resin or the air, allow these photo-sensors to become themselves optimal detectors for charged particle identification both via time of flight or via photon counting measurements.
@@ Acknowledgments
The present work is largely the result of the efforts and ingenuity of Andrea Alici, who unfortunately passed away just a few days before its publication. His memory will live on among his colleagues.

### v2

@@ 1 Introduction
In [3, 6, 5, 7] it was quantitatively shown that the Cherenkov radiation produced at the passage of a charged particle through the standard, factory-supplied protection layer of the SiPM itself can produce very high number of firing cells (SPADs with area tens of microns square). Considering a refractive index of 1.56 and a thickness of few hundreds micron for the Silicone material of the protection layer, the Cherenkov angle reaches $\approx$50 degrees, covering a quite large area of few mm2. In these conditions it is possible to directly detect charged particles and reach very good efficiency and time resolution around 20 ps [5, 7]. These results opened the possibility to use just a single SiPM as an optimal detector for Time Of Flight (TOF) measurements.
The results reported in the above papers also indicate a very reduced number of firing SPADs when no protection layer is present. For such sensors, events with more than one SPAD are mainly due to cross-talk processes [5, 7].
@@
To test the above possibility, in this paper data collected at T10 test beam at CERN with momenta from 1.5 to 10 GeV/$c$ are used. The apparatus is the same described in [5]. At lower momenta the fraction of electrons/positrons in the beam increases thus allowing to compare the SiPM (without protection) signal with the one observed for pions and protons. In Section 2 the experimental set-up and the analysis methods are described. The results are reported in Section 3 and will be compared with a simple simulation in Section 4. In Section 5 a possible optimization of the parameters involved in such study (length of air traversed, different traversed gas, different SiPM technology) is discussed.
@@ 2.2 Beam test setup
The SiPMs were tested at the CERN PS T10 particle beam in the momentum range 1.5-10 GeV/$c$. The beam composition, evaluated via simulation and data, is reported in fractions with few % uncertainties, in [14]: while at high momenta the beam is mainly made of protons and $\pi^{+}$, at lower energies the fraction of positrons increases and becomes dominant. In the following, we will neglect the charge, which is uninfluential for our purposes, and use generically the terms pions and electrons.
@@ 2.3 Analysis method
The analyzed data were collected with a beam momentum of 1.5 GeV/$c$ (a second sample at 10 GeV/c was also collected to check the simulation). The reason for this choice is twofold. First, it enhances the electron component in the beam, which is composed of approximately 60% electrons, 28% pions, and 12% protons [14]. Notice that these fractions are estimated with few % uncertainties and such variations do not affect the results reported in Section 3. Second, it enables the separation of electrons and pions from protons by exploiting the time of flight between the signals from the two LGAD detectors placed at the ends of the telescope, as visible in Fig. 1 left (more details in [5]). In the figure a q-Gaussian fit [13] was applied on both distributions.
Figure 1: Left panel: time of flight of the beam particles computed as the time difference between the two LGADs in the telescope at 1.5 GeV/$c$. The fits are performed with two q-Gaussian, the right one relative to the proton content and the left one relative to pions/electrons. Right panel: LGAD time sum versus time difference distribution with the box cut for protons (red) and electron/pions (green).
At this beam momentum, electrons are the only particles that produce Cherenkov radiation in the air gap between sensors in the telescope. Electrons and pions cannot be distinguished using the LGADs’ time information; however, their charge spectra can still be determined. The proton charge spectrum is used as a proxy for pions normalized to the expected pion fraction [14], since neither pions nor protons are able to generate Cherenkov radiation in air.
The protons sample has been selected with rectangular cuts on the LGAD time sum versus time difference distribution (Fig. 1 right) as $0.2<t_{LGAD1}-t_{LGAD2}<0.4$ ns and $1.2<t_{LGAD1}+t_{LGAD2}<-0.85$ ns, while for electrons and pions the cuts are $0<t_{LGAD1}-t_{LGAD2}<0.16$ ns and $1.2<t_{LGAD1}+t_{LGAD2}<-0.7$ ns. With these cuts the fraction of protons remaining in the electron/pion sample is below 1%, while the electron/pion contamination in the proton sample is $\approx$ 4-5%.
@@ 3 Results
The measured amplitudes for the proton sample is reported in Fig. 2 top-left: as expected, it is peaked at one SPAD ($\sim$31 mV), with the entries at 0 mV (0 SPAD) compatible with the value of the fill factor of the sensor and with peaks at 2 and 3 SPADs due to crosstalk and possible electron contamination ( see 2.3 ). The amplitude distribution for pions is expected to be identical to that of protons. The electron distribution is then obtained by taking the amplitude distribution of events identified as electron/pion candidates and subtracting the proton distribution, normalized to the expected pion fraction of 32%[14]. The pion/electron spectrum is reported in Fig. 2 top-right with the estimated fraction of pions and electrons. The electron spectrum, after the pion subtraction, is instead reported in Fig. 2 bottom.
@@ 4 Comparison with simulation
The results reported in the previous section rely mainly on the purity of the selected proton sample and on an accurate knowledge of the actual composition of the particle beam. To estimate the impact of these experimental uncertainties the results were cross-checked against the predictions of a toy Monte Carlo (MC) simulation to assess their robustness. The simulation works as follows: two hits are randomly generated on the two LGADs and the track is then propagated to the SiPM. The number of Cherenkov photons created in the air volume upstream the sensor has been calculated based on the Frank-Tamm formula [8]. The detection efficiency for the charged particle was assumed equal to the fill factor according to the manufacturer’s specifications, as well as the crosstalk and the PDE as a function of the photon wavelength and SiPM overvoltage. The amplitude distribution is obtained by multiplying the number of fired cells by the average amplitude of 1 SPAD signal assumed as 31 mV according to the acquired data. A gaussian spread of 3.3 mV was applied driven by data observation. Given the binary nature of the SPAD response, multiple photons hitting the same pixel are counted as one. A still-not-fully-satisfactory modeling of the noise is applied to the simulated data based on the measured RMS of the baseline. However, no explicit simulation of the electronic jitter or of the detailed baseline fluctuations described in Sec. 2.3 is included, which may affect in particular the population of events between peaks and the 0-signal region.
@@
A qualitative comparison between the measured spectra and MC predictions is shown in Fig. 3. In the left pad, the measured amplitude spectrum for protons with a momentum of 1.5 GeV/$c$ is compared with simulations. Here and in the following, the two histograms under comparison are normalized to the same number of entries. The measured spectrum shows a small peak at 3 SPADs which is not predicted by the toy MC. This discrepancy ($\sim 3\%$ of events with more than 2 SPADs) mainly arise from a small electron contamination in the proton sample selected by LGAD timing information, $\sim 2-3\%$; the remaining part may originate from events in which crosstalk affects more than one pixel, which is not included in the MC, or from particle near the border of two SPADs. In the same figure the right panel shows a comparison between measured amplitude spectra and simulations at a beam momentum of 10 GeV/$c$. At this momentum, the beam consists mainly of pions (30%) and protons (70%), with pions producing Cherenkov radiation in air, while protons remain below threshold [^3] As from [14], the fractions used in this comparison are a simplification and the uncertainties of few % may affect especially the null signal . In this case, the agreement between the two distributions is overall satisfactory in terms of peak positions and relative populations, although discrepancies in the inter-peak and 0-signal regions are still visible. Finally, Fig. 4 compares the measured and simulated amplitude spectra for electrons at 1.5 GeV/$c$ : here too, the comparison shows good agreement.
It should be stressed that this toy MC comparison is intended to provide a qualitative support to the interpretation of the experimental observations. A more detailed simulation, including a refined description of the electronic noise and baseline response, would be required for a fully quantitative assessment.
@@ 5 A simple optimization of the key parameters
Before using the toy MC to obtain an approximate estimate of the potential performance of such a system, a number of parameters were optimized. It was observed that the SiPM coverage has a significant impact, with larger covered areas able to collect more photons produced by the incident charged particle. No significant changes in performance were observed reducing the cell size, within the range tested. A comparison was therefore made between SiPMs with active areas of 3.20 $\times$ 3.12 mm2, as the one used in the test beam data, and a hypothetical 6 $\times$ 6 mm2, both featuring square SPADs with a pitch of 40 µm. To determine the optimal thickness of the air layer in front of the sensor, the simulation considers particles impinging the SiPM in the center. The over-voltage was set to 6 V to profit of the improved PDE.
@@
The optimization was carried out using the plot shown in Fig. 5, where the electron efficiency was calculated considering particles with 1.5 GeV/$c$ of momentum and by selecting only events in which three or more SPADs were fired. As can be seen, while below 4 cm the curves relative to the two different size SiPM overlap, at higher distances the larger-area SiPM exhibits a significantly higher efficiency and this is due to the larger surface area, which allows more Cherenkov photons to be collected.
@@
Figure 5: Electron detection efficiency versus the distance in air traversed by the particle. Electrons of 1.5 GeV/$c$ and two different SiPMs surfaces are considered. The toy Monte Carlo parameters used are: NUV-HD-MT, OV = 6 V, Thr = 3 SPAD.
The figure also indicates that for the 6 $\times$ 6 mm2 sensor beyond 15 cm of air no appreciable variations in the number of collected photons are observed due to the limited size of the simulated sensor. In the following simulations, the response of the larger SiPM will be considered using a distance of 15 cm of traversed air. Notice that the goal of this paper is to investigate the method to identify particles via the SiPM response and there is no attempt to optimize the detector configuration as it should be done for specific applications.
Since electron identification is performed by considering signals with an amplitude higher than the cross-talk value, it is clear that this value represents a fundamental parameter. Recently, the FBK NUV-HD SiPM technology has been upgraded with metal-filled deep trench isolation (NUV-HD-MT [12]) between the SiPM cells, which enables an almost complete suppression of internal optical cross-talk. In the simulation, therefore, in order to select the optimal configuration for the proposed detector, the cross-talk values measured by FBK for this new technology were used. The PDE also impacts performance, as it parameterizes the probability of detecting the produced Cherenkov photons. Assuming the validity of the simulation also at overvoltages higher than 2 V, a value of 6 V was chosen to profit of the higher PDE ( from about 40% to 60%). The cross-talk and PDE values corresponding to OV = 6 V were taken from [12]: possible uncertainties related to the extrapolation from 2 to 6 V overvoltage would require a careful check with experimental data in future beam test campaigns.
@@
Starting from the excellent timing capability of SiPM with a protection layer reported in [5], the approach discussed in the present paper suggests that configurations with two consecutive SiPMs, one without and one with protection layer, with an air gap in front of the first one, could be used to simultaneously support electron identification and TOF measurements, potentially covering PID and particle rejection in different momentum ranges.
@@ 6 Conclusions
Together with the already published results on timing performances of SiPMs with a protection layer [5, 7], the present results underline that the Cherenkov light produced in the simplest radiators in front of the SiPM, i.e. the factory-supplied silicone protective resin or the air, allow these photo-sensors to become themselves optimal detectors for charged particle identification both via time of flight or via photon counting measurements.
@@ Acknowledgments
The present work is mainly the result of the efforts and ingenuity of Andrea Alici, who unfortunately passed away just a few days before its publication. His memory will live on among his colleagues.